# DynaSpec: Context-aware Dynamic Speculative Sampling for Large-Vocabulary Language Models

### v2

@@ Abstract
1 Introduction
Over the past few years, large language models (LLMs) (brown2020language; touvron2023llama; liu2024deepseek; jiang2024mixtral) have improved dramatically, reshaping products and workflows across sectors. Beyond algorithmic advances, much of this progress has come from scaling data and especially model parameters (kaplan2020scaling; hoffmann2022training). Yet these gains come with a cost: inference becomes compute-heavy and latency-sensitive, which limits real-time applications (kwon2023efficient; yuan2024llm). As agentic systems (yao2023react; wang2024survey) and reasoning-heavy workflows (e.g., chain-of-thought) (wei2022chain; zhang2024chain) proliferate, the demand for faster inference keeps rising. Speculative decoding (chen2023accelerating; leviathan2023fast; li2024eagle; li2024eagle2) is a practical way to accelerate generation. A smaller draft model proposes multiple tokens so the target model need not run at every step, while the rejection sampling procedure guarantees that the final output distribution matches that of the target model. In the worst case, quality and latency revert to standard decoding; in typical cases, throughput improves substantially.
Recently scaling laws for vocabulary (tao2024scaling; huangover; yu2025scaling) suggests that larger models benefit from larger vocabularies, and deployed LLMs have indeed expanded their vocabularies over time. For example, tokenizer vocabularies have scaled from 32k entries in Llama-2 (touvron2023llama) to 128k in Llama-3 (grattafiori2024llama), 129k in DeepSeek-V3 (liu2024deepseek), 152k in Qwen-2.5 (yang2025qwen3), and 262k in Gemma-3 (team2025gemma). While the large vocabulary size contributes only marginally to the inference cost of the target model’s output layers given its overall model size, it imposes a substantial computational burden on the draft model during speculative decoding, as the draft model is much smaller and the output layer constitutes a larger fraction of its total computation. thereby diminishing the overall pipeline efficiency. To mitigate this, FR-Spec (zhao-etal-2025-fr) and Vocabtrim (goel2025vocabtrim) replace the full vocabulary with a fixed shortlist—typically the top-p% of the most frequent tokens measured on a reference corpus. This speeds up drafting, but it is suboptimal for two reasons: (1) frequency lists are corpus-dependent and often fail to generalize across benchmarks, requiring substantial retuning; and (2) a static subset of the token vocabulary can degrade performance on tasks that rely on rare or domain-specific tokens, where maintaining diversity in the candidate set is critical for high-quality outputs.
In this work, we move beyond fixed shortlists and introduce a dynamic vocabulary head. We first define meta-labels as coarse token clusters: the vocabulary is partitioned into $M\!\ll\!|V|$ clusters $\{C_{m}\}_{m=1}^{M}$ by clustering column-normalized LM-head weights. Conditioned on embedding and features, a lightweight meta-classifier scores clusters and we evaluate the drafter’s LM head only on the union of the selected clusters $\mathcal{V}_{S}=\bigcup_{m\in\mathcal{K}}\!C_{m}$, while verification still runs over the full vocabulary, preserving exactness. This coarse-to-fine routing, inspired by jiang2021lightxml and  kharbanda2022cascadexml, replaces an $\mathcal{O}(|V|d)$ head with $\mathcal{O}\!\big((M+|\mathcal{V}_{S}|)d\big)$. Making the support context-dependent leads to a consistent increase (across datasets) in the mean accepted length relative to the static subset proposed in zhao-etal-2025-fr. However, a dynamic (indexed) head introduces gathered matmul overhead and can be slower than a fixed, pre-trimmed dense head. We therefore couple the dynamic head with a position-aware cluster budget: early tokens receive larger shortlists, and the budget decays with position. As argued in recent works, li2025gumiho; li2024eagle2, prioritizing earlier tokens improves expected acceptance, while shrinking the shortlist later reduces latency, yielding net end-to-end gains without altering verification. Finally, we show the same position-aware schedule can retrofit frequency-ranked shortlists (FR-Spec) (zhao-etal-2025-fr); see Appendix A.1 for details.
Contributions.
To reconcile acceptance gains from a dynamic head with the matmul overhead it introduces, DynaSpec combines context-aware routing with a position-aware budget. Concretely: (i) we introduce meta-labels (coarse token clusters) and a lightweight meta-classifier that selects a context-dependent candidate shortlist of clusters on which the drafter operates, while verification remains exact on the full vocabulary; (ii) we show that dynamic, context-conditioned supports dominate any static subset in expected acceptance; (iii) we mitigate indexed-head latency via a position-aware schedule $k_{\mathrm{c}}(t)$, allocating more candidate clusters early and fewer later, with a fused index+GEMM kernel, yielding lower draft-time compute without sacrificing early-step acceptance; and (iv) empirically, across seven tasks and LLaMA models with $128\mathrm{k}$ vocabularies, DynaSpec increases mean accepted length over fixed-shortlist baselines. The approach is plug-in compatible with EAGLE-style pipelines and orthogonal to system-side optimizations, making it a practical drop-in for large-vocabulary deployments.
@@ 2.1 LLMs with Large Vocabulary
As LLMs have grown in scale, so too have their vocabularies. Early models such as GPT-3 (brown2020language) and LLaMA-2 (touvron2023llama) used vocabularies of $32\mathrm{k}$ tokens, while more recent releases exceed $100\mathrm{k}$ (e.g., $128\mathrm{k}$ in LLaMA-3 (grattafiori2024llama), $152\mathrm{k}$ in Qwen-2.5 (yang2025qwen3), and $262\mathrm{k}$ in Gemma-3 (team2025gemma)). Large vocabularies are especially common in multilingual and high-coverage LLMs: mT5 (xue2020mt5) adopts a $250\mathrm{k}$ SentencePiece (kudo2018sentencepiece) vocabulary to improve cross-lingual coverage, and PaLM uses a $256\mathrm{k}$ unigram vocabulary to shorten sequences and better encode code and symbols (chowdhery2023palm). Recently, empirical scaling laws suggest that larger vocabularies improve expressivity and perplexity (tao2024scaling; yu2025scaling), but they also magnify inference cost: the output projection grows linearly with vocabulary size, which can dominate latency in small drafters.
@@ 2.2 Speculative Decoding
In line with Speculative Decoding, early work accelerates greedy generation (stern2018blockwise; sun2021instantaneous), and subsequent work extends the idea to non-greedy sampling with distributional guarantees (leviathan2023fast; chen2023accelerating; xia2022speculative). These methods differ primarily in how they form proposals: some reuse information already present in the prompt via retrieval-guided drafting—effective when outputs strongly overlap inputs (saxena2023prompt; yang2023inference; he2024rest)—whereas others learn parametric drafters for general-purpose speedups, ranging from parallel MLP heads (and serial variants) to lightweight transformer blocks that condition on richer context (cai2024medusa; ankner2024hydra). The EAGLE series further refines this paradigm: EAGLE employs lightweight transformer drafters, while EAGLE-2 introduces dynamic draft trees to boost efficiency (li2024eagle; li2024eagle2). Orthogonal advances in training and systems—such as adaptive exploration, cache reuse, and efficient serving stacks—increase accepted tokens per verification and reduce overhead (kwon2023efficient; zheng2024sglang; dao2023flashattention). Complementing these directions, MagicDec analyzes the latency–throughput trade-off and shows that speculative decoding remains effective at large batch sizes and with long contexts, whereas TriForce targets long-context regimes by compressing the drafter’s KV cache and coordinating reuse across draft steps (sadhukhan2024magicdec; sun2024triforce). Most recently, timor2025accelerating proposes lossless speculative decoding for heterogeneous vocabularies, relaxing the same-tokenizer requirement between drafter and target; meanwhile, li2025gumiho combines Medusa-style heads (cai2024medusa) with the EAGLE (li2024eagle2) framework. We target a different bottleneck: as vocabularies expand, the drafter’s output projection increasingly dominates latency.
@@ 2.3 Accelerating Large Vocabularies
The computational bottleneck of large output spaces or vocabularies has been recognized in both language modeling and neighboring fields. Within LLMs, FR-Spec (zhao-etal-2025-fr) and VocabTrim (goel2025vocabtrim) mitigate drafter overhead by pruning the output layer to a fixed shortlist of high-frequency tokens. While effective in reducing latency, these static subsets may suppress rare or domain-specific tokens, lowering acceptance in speculative decoding. Other work on large vocabulary LLM training, such as CCE (wijmans2024cut), reduces peak memory by fusing cross-entropy directly into the kernel, avoiding explicit logit materialization. Our approach is motivated by research in extreme classification, where the goal is to accelerate learning and inference over output spaces with millions of labels. Methods such as LightXML (jiang2021lightxml), CascadeXML (kharbanda2022cascadexml), and related meta classifier based strategies (zhang2021fast; kharbanda2023inceptionxml) adopt coarse-to-fine or cluster-based mechanisms that limit computation to a small subset of relevant labels per input. These ideas directly inspire DynaSpec: rather than evaluating the drafter across the entire vocabulary or a static shortlist, we employ a lightweight meta-classifier to select a small number of context-relevant clusters, retaining the coverage of rare tokens while reducing compute.
3 Theoretical Analysis
Let $T_{T}$ denote the per-token latency of the target model under standard autoregressive decoding. In speculative decoding, let $T_{D}$ be the draft model’s per-token generation time, and let $T_{V}(\gamma)$ be the time for the target to verify $\gamma$ proposed tokens. Given $\alpha\in[0,1]$ as the expected acceptance rate and $\gamma\in\mathbb{N}$ the speculation length, following standard derivations (leviathan2023fast), the expected number of tokens produced per verification step is
@@ 3 Theoretical Analysis
$T_{\mathrm{SD}}=\frac{\gamma\cdot T_{D}+T_{V}(\gamma)}{\Omega(\alpha,\gamma)}.$
@@
$\mathrm{Speedup}=\frac{T_{T}}{T_{\mathrm{SD}}}=\frac{T_{T}\cdot\Omega(\alpha,\gamma)}{\gamma\cdot T_{D}+T_{V}(\gamma)}.$ (2)
3.1 Large-Vocabulary Bottleneck on Speedup
@@
As shown in Equation 2 the speedup of speculative decoding is determined by the verification time $T_{V}(\gamma)$, the draft cost $T_{D}$, and the expected generation length $\Omega(\alpha,\gamma)$. As the tokenizer vocabulary grows, the target-side terms $T_{T}$ and $T_{V}(\gamma)$ are comparatively insensitive to $|V|$, since their per-token runtime is dominated by stacked attention and MLP compute across many layers. In contrast, the drafting time per-token decomposes into a transformer core term $\Theta(L_{D}d^{2}+L_{D}Sd)$ (attention and MLP) and a vocabulary-head term $\Theta(|V|d)$. As $V$ increases, the $\Theta(|V|d)$ component can overtake the core, inflating $T_{D}$ and—because it is multiplied by $\gamma$ in the speedup denominator—either reducing the achievable speedup or forcing a smaller $\gamma$.
@@
Restricting the draft to a smaller subset of the vocabulary effectively reduces the $\Theta(|V|d)$ term and can lower $T_{D}$. However, if this subset is not context-aware, the acceptance rate $\alpha$ may drop, shrinking $\Omega(\alpha,\gamma)$ and hence offsetting the latency gains, as we discuss next.
3.2 Impact of Static Sub-Vocabulary Constraints on Acceptance
 [^1] For details, follow section 3 in (leviathan2023fast)
We use context $c$ interchangeably with the conditioning prefix $x_{<t}$. Assuming $p_{c}(x)\;\equiv\;p(x_{t}{=}x\mid x_{<t}{=}c)$, the target distribution and $q_{c}(x)\;\equiv\;q(x_{t}{=}x\mid x_{<t}{=}c),$ for the draft model over the full vocabulary $V$. The acceptance probability (acceptance rate) is
$\beta(c;q_{c})$ $=\;\sum_{x\in V}\min(p_{c}(x),\,q_{c}(x))$ (3)
$=\;\mathbb{E}_{x\sim q_{c}}\!\left[\min\!\Big(1,\tfrac{p_{c}(x)}{q_{c}(x)}\Big)\right].$ (4)
Under the standard i.i.d. assumption for $\beta$ , the average acceptance is
$\alpha\;:=\;\mathbb{E}_{c}\!\left[\beta(c;q_{c})\right].$ (5)
Lemma 1 (Acceptance on a fixed support). For any context $c$ and any subset $S\subseteq V$ with $|S|=k$, following Equation 3, the acceptance rate can be written as
$A^{\!*}(c;S)$ $:=\;\max_{q_{c}:\ \mathrm{supp}(q_{c})\subseteq S}\ \sum_{x\in V}\min\{p_{c}(x),q_{c}(x)\}$
$=\;\sum_{x\in S}p_{c}(x)\,.$
That is, $A^{\!*}(c;S)$ is the maximum of $\beta(c;q_{c})$ over all $q_{c}$ whose support is contained in $S$.
Proof. For $x\notin S$, the overlap $\min\{p_{c}(x),q_{c}(x)\}=0$. For $x\in S$, $\min\{p_{c}(x),q_{c}(x)\}\leq p_{c}(x)$ with equality iff $q_{c}(x)\geq p_{c}(x)$. Normalizing $p_{c}$ on $S$ makes this hold for all $x\in S$. Summing over $x\in S$ gives the value. $\square$
Lemma 2 (Oracle or Pointwise optimal subset). Let $p_{c}^{\downarrow}(1)\geq\cdots\geq p_{c}^{\downarrow}(|V|)$ be the sorted probabilities and define the top-$k$ mass $H_{k}(p_{c}):=\sum_{i=1}^{k}p_{c}^{\downarrow}(i)$. Then
$\max_{|S|=k}A^{\!*}(c;S)\;=\;H_{k}(p_{c}),$
attained by any top-$k$ subset under $p_{c}$ (ties allowed).
Proof. By Lemma 1, $\max_{|S|=k}A^{\!*}(c;S)=\max_{|S|=k}\sum_{x\in S}p_{c}(x)=H_{k}(p_{c})$ (rearrangement/majorization). $\square$
Theorem A (Oracle dynamic dominates any static subset). Let $S_{c}^{\star}$ be the oracle top-$k$ set for context $c$, i.e., any measurable top-$k$ subset under $p_{c}$. For any static baseline $S_{\mathrm{stat}}$ with $|S_{\mathrm{stat}}|=k$,
$\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{c}^{\star})\big]=\;\mathbb{E}_{c}\!\big[H_{k}(p_{c})\big]\;\geq\;\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{\mathrm{stat}})\big]$
$=\;\mathbb{E}_{c}\!\Big[\sum_{x\in S_{\mathrm{stat}}}p_{c}(x)\Big],$
with strict inequality whenever, on a set of contexts of positive probability, the (unique) top-$k$ set differs from $S_{\mathrm{stat}}$.
Proof. Pointwise, $A^{\!*}(c;S_{c}^{\star})=H_{k}(p_{c})\geq\sum_{x\in S_{\mathrm{stat}}}p_{c}(x)=A^{\!*}(c;S_{\mathrm{stat}})$ by Lemma 2. Take expectations; strictness follows from a strict pointwise gap on a set of positive measure. $\square$
Consequence (expected tokens per verification). Fix speculation length $\gamma$. Since $\Omega(\alpha,\gamma)$ is strictly increasing in $\alpha$, Theorem A implies
$\Omega\!\big(\alpha_{\mathrm{dyn}}^{\star},\gamma\big)\;\geq\;\Omega\!\big(\alpha_{\mathrm{stat}},\gamma\big),$
where $\alpha_{\mathrm{dyn}}^{\star}$ denotes the optimal expected acceptance rate. A dynamic context-aware subset yields at least as many, and typically more generated tokens per verification step by retaining more target probability mass within the drafting support than any static subset.
In the next section, we propose DynaSpec, a context-aware, meta-classifier–driven dynamic vocabulary head that reduces the drafter’s per-token cost $T_{D}$ while maintaining a high $\Omega(\alpha,\gamma)$; verification continues over the full vocabulary, preserving exactness.
4 DynaSpec
Unlike a fixed output head, DynaSpec selects tokens dynamically based on context. It uses a lightweight meta-classifier (the router) to route hidden state of draft model to coarse token clusters, decoding only over the resulting shortlist of tokens within the clusters that are selected by the meta-classifier.
4.1 Preliminary

### v3

@@ Abstract
Speculative decoding accelerates LLM inference by letting a small drafter propose multiple tokens which a large target model verifies once per speculation step. As vocabularies scale past $10^{5}$ tokens, verification cost in the target model is largely unchanged, but the drafter can become bottlenecked by its $\mathcal{O}(|V|d)$ output projection. Recent approaches (e.g., FR-Spec, VocabTrim) mitigate this by restricting drafting to a fixed, frequency-ranked shortlist; however, such static truncation is corpus-dependent and suppresses rare or domain-specific tokens, reducing acceptance and limiting speedups. We propose DynaSpec, a context-dependent dynamic shortlisting mechanism for large-vocabulary speculative decoding. DynaSpec trains lightweight meta-classifiers that route each context to a small set of coarse token clusters; the union of the top-selected clusters defines the drafter’s shortlist, while the target model still verifies over the full vocabulary, preserving exactness. Systems-wise, routing is overlapped with draft computation via parallel execution streams, reducing end-to-end overhead. Across standard speculative decoding benchmarks, DynaSpec consistently improves mean accepted length—recovering 98.4% of full-vocabulary performance for Llama-3-8B versus 93.6% for fixed-shortlist baselines—and achieves up to a 2.23× throughput gain compared to 1.91× for static approaches on the dataset with rare tokens.
1 Introduction
Over the past few years, large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Liu et al., 2024; Jiang et al., 2024) have improved dramatically, reshaping products and workflows across sectors. Beyond algorithmic advances, much of this progress has come from scaling data and especially model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Yet these gains come with a cost: inference becomes compute-heavy and latency-sensitive, which limits real-time applications (Kwon et al., 2023; Yuan et al., 2024). As agentic systems (Yao et al., 2023; Wang et al., 2024) and reasoning-heavy workflows (e.g., chain-of-thought) (Wei et al., 2022; Zhang et al., 2024) proliferate, the demand for faster inference keeps rising. Speculative decoding (Chen et al., 2023; Leviathan et al., 2023; Li et al., 2024b, a) is a practical way to accelerate generation. A smaller draft model proposes multiple tokens so the target model need not run at every step, while the rejection sampling procedure guarantees that the final output distribution matches that of the target model. In the worst case, quality and latency revert to standard decoding; in typical cases, throughput improves substantially.
Recently scaling laws for vocabulary (Tao et al., 2024; Huang et al., ; Yu et al., 2025) suggest that larger models benefit from larger vocabularies, and deployed LLMs have indeed expanded their vocabularies over time. For example, tokenizer vocabularies have scaled from 32k entries in Llama-2 (Touvron et al., 2023) to 128k in Llama-3 (Grattafiori et al., 2024), 129k in DeepSeek-V3 (Liu et al., 2024), 152k in Qwen-2.5 (Yang et al., 2025), and 262k in Gemma-3 (Team et al., 2025). While the large vocabulary size contributes only marginally to the inference cost of the target model’s output layers given its overall model size, it imposes a substantial computational burden on the draft model during speculative decoding, as the draft model is much smaller and the output layer constitutes a larger fraction of its total computation, thereby diminishing the overall pipeline efficiency. To mitigate this, FR-Spec (Zhao et al., 2025) and Vocabtrim (Goel et al., 2025) replace the full vocabulary with a fixed shortlist—typically the top-p% of the most frequent tokens measured on a reference corpus. This speeds up drafting, but it is suboptimal for two reasons: (1) frequency lists are corpus-dependent and often fail to generalize across benchmarks, requiring substantial retuning; and (2) a static subset of the token vocabulary can degrade performance on tasks that rely on rare or domain-specific tokens, where maintaining diversity in the candidate set is critical for high-quality outputs.
In this work, we move beyond fixed shortlists and introduce a dynamic vocabulary head. We first define meta-labels as coarse token clusters: the vocabulary $V$ is partitioned into $M\!\ll\!|V|$ clusters $\{C_{m}\}_{m=1}^{M}$ by clustering column-normalized LM-head weights. Conditioned on current context (in the form of embedding and features), a lightweight meta-classifier scores clusters and we evaluate the drafter’s LM head only on the union of the selected clusters $V_{S}=\bigcup_{m\in\mathcal{K}}\!C_{m}$, while verification still runs over the full vocabulary, preserving exactness. This coarse-to-fine routing, inspired by Jiang et al. (2021) and  Kharbanda et al. (2022), replaces an $\mathcal{O}(|V|d)$ head with $\mathcal{O}\!\big((M+|V_{S}|)d\big)$. Making the support context-dependent leads to a consistent increase (across datasets) in the mean accepted length relative to the static subset proposed in Zhao et al. (2025). However, a dynamic (indexed) head introduces gathered matmul overhead and can be slower than a fixed, pre-trimmed dense head. We, therefore, couple the dynamic head with a position-aware cluster budget: early tokens receive larger shortlists, and the budget decays with position. As argued in recent works, Li et al. (2025, 2024a), prioritizing earlier tokens improves expected acceptance, while shrinking the shortlist later reduces latency, yielding net end-to-end gains without altering verification. Finally, we show the same position-aware schedule can retrofit frequency-ranked shortlists (FR-Spec) (Zhao et al., 2025).
Contributions
To address the drafter bottleneck in large-vocabulary speculative decoding, we introduce DynaSpec, a context-aware dynamic vocabulary head that trades a full $\mathcal{O}(|V|d)$ projection for a routed, sparse head while preserving exact verification. Specifically:
  • Context-aware dynamic shortlisting via cluster routing. We partition the vocabulary into coarse clusters and train a lightweight router to select a small set of clusters per decoding step; the drafter evaluates logits only on the union of selected clusters, while the target model continues to verify on the full vocabulary.
  • Theory: context-conditioned supports dominate static truncation. We provide an acceptance analysis showing that, for a fixed budget, context-dependent supports can strictly improve expected overlap/acceptance relative to any fixed subset, and we relate router quality to an explicit “routing regret” term.
  • Systems: amortizing routed-head overhead. To reconcile dynamic shortlisting with the overhead of indexed (gathered) matmul, we (i) run routing on a parallel CUDA stream so its latency is largely hidden under draft computation, and (ii) introduce a position-aware cluster budget that allocates larger shortlists to early draft steps and decays thereafter; we further fuse index selection with the shortlisted-head GEMM in a custom kernel.
@@ 2.1 LLMs with Large Vocabulary
As LLMs have grown in scale, so too have their vocabularies. Early models such as GPT-3 (Brown et al., 2020) and LLaMA-2 (Touvron et al., 2023) used vocabularies of $32\mathrm{k}$ tokens, while more recent releases exceed $100\mathrm{k}$ (e.g., $128\mathrm{k}$ in LLaMA-3 (Grattafiori et al., 2024), $152\mathrm{k}$ in Qwen-2.5 (Yang et al., 2025), and $262\mathrm{k}$ in Gemma-3 (Team et al., 2025)). Large vocabularies are especially common in multilingual and high-coverage LLMs: mT5 (Xue et al., 2020) adopts a $250\mathrm{k}$ SentencePiece (Kudo and Richardson, 2018) vocabulary to improve cross-lingual coverage, and PaLM uses a $256\mathrm{k}$ unigram vocabulary to shorten sequences and better encode code and symbols (Chowdhery et al., 2023). Recently, empirical scaling laws suggest that larger vocabularies improve expressivity and perplexity (Tao et al., 2024; Yu et al., 2025), but they also magnify inference cost: the output projection grows linearly with vocabulary size, which can dominate latency in small drafters.
@@ 2.2 Speculative Decoding
In line with Speculative Decoding, early work accelerates greedy generation (Stern et al., 2018; Sun et al., 2021), and subsequent work extends the idea to non-greedy sampling with distributional guarantees (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2022). These methods differ primarily in how they form proposals: some reuse information already present in the prompt via retrieval-guided drafting—effective when outputs strongly overlap inputs (Saxena, 2023; Yang et al., 2023; He et al., 2024)—whereas others learn parametric drafters for general-purpose speedups, ranging from parallel MLP heads (and serial variants) to lightweight transformer blocks that condition on richer context (Cai et al., 2024; Ankner et al., 2024). The EAGLE series further refines this paradigm: EAGLE employs lightweight transformer drafters, while EAGLE-2 introduces dynamic draft trees to boost efficiency (Li et al., 2024b, a). Orthogonal advances in training and systems—such as adaptive exploration, cache reuse, and efficient serving stacks—increase accepted tokens per verification and reduce overhead (Kwon et al., 2023; Zheng et al., 2024; Dao, 2023). Complementing these directions, MagicDec analyzes the latency–throughput trade-off and shows that speculative decoding remains effective at large batch sizes and with long contexts, whereas TriForce targets long-context regimes by compressing the drafter’s KV cache and coordinating reuse across draft steps (Sadhukhan et al., 2024; Sun et al., 2024). Most recently, Timor et al. (2025) proposes lossless speculative decoding for heterogeneous vocabularies, relaxing the same-tokenizer requirement between drafter and target; meanwhile, Li et al. (2025) combines Medusa-style heads (Cai et al., 2024) with the EAGLE (Li et al., 2024a) framework. We target a different bottleneck: as vocabularies expand, the drafter’s output projection increasingly dominates latency.
@@ 2.3 Accelerating Large Vocabularies
The computational bottleneck of large output spaces or vocabularies has been recognized in both language modeling and neighboring fields. Within LLMs, FR-Spec (Zhao et al., 2025) and VocabTrim (Goel et al., 2025) mitigate drafter overhead by pruning the output layer to a fixed shortlist of high-frequency tokens. While effective in reducing latency, these static subsets may suppress rare or domain-specific tokens, lowering acceptance in speculative decoding. Other work on large vocabulary LLM training, such as CCE (Wijmans et al., 2024), reduces peak memory by fusing cross-entropy directly into the kernel, avoiding explicit logit materialization. Our approach is motivated by research in extreme classification, where the goal is to accelerate learning and inference over output spaces with millions of labels. Methods such as LightXML (Jiang et al., 2021), CascadeXML (Kharbanda et al., 2022), and related meta classifier based strategies (Zhang et al., 2021; Kharbanda et al., 2023) adopt coarse-to-fine or cluster-based mechanisms that limit computation to a small subset of relevant labels per input. These ideas directly inspire DynaSpec: rather than evaluating the drafter across the entire vocabulary or a static shortlist, we employ a lightweight meta-classifier to select a small number of context-relevant clusters, retaining the coverage of rare tokens while reducing compute.
3 Computational bottleneck of speculative decoding in Large-vocabulary LMs
Under standard autoregressive decoding, let $\tau_{T}$ denote the per-token latency of the target model. With speculative decoding, let $\tau_{D}$ be the draft model’s per-token generation time, and $\tau_{V}(\gamma)$ be the time for the target to verify $\gamma$ the proposed tokens. Given $\alpha\in[0,1]$ as the expected acceptance rate and $\gamma\in\mathbb{N}$ the speculation length, the expected number of tokens produced per verification step is (per equation (1) in Leviathan et al. (2023)) :
@@ 3 Computational bottleneck of speculative decoding in Large-vocabulary LMs
$\tau_{\mathrm{SD}}=\frac{\gamma\cdot\tau_{D}+\tau_{V}(\gamma)}{\Omega(\alpha,\gamma)}.$
@@
$\mathrm{Speedup}=\frac{\tau_{T}}{\tau_{\mathrm{SD}}}=\frac{\tau_{T}\cdot\Omega(\alpha,\gamma)}{\gamma\cdot\tau_{D}+\tau_{V}(\gamma)}.$ (2)
@@
As shown in Equation 2 the speedup of speculative decoding is determined by the verification time $\tau_{V}(\gamma)$, the draft cost $\tau_{D}$, and the expected generation length $\Omega(\alpha,\gamma)$. Since the target LMs main computation during inference is dominated by multiple stacked attention and MLP layers, the terms $\tau_{T}$ and $\tau_{V}(\gamma)$ remain relatively unaffected by increasing vocabularies. On the other hand, the drafting time per-token decomposes into a single transformer layer (attention and MLP) with core term $\Theta(L_{D}d^{2}+L_{D}Sd)$ and a vocabulary-head term $\Theta(|V|d)$. As $V$ increases, the $\Theta(|V|d)$ component can overtake the core, inflating $\tau_{D}$ and—because it is multiplied by $\gamma$ in the speedup denominator—either reducing the achievable speedup or forcing a smaller $\gamma$.
@@
Restricting the draft to a smaller subset of the vocabulary, a crude approach taken in FR-Spec Zhao et al. (2025) and Vocab-Trim Goel et al. (2025), effectively reduces the $\Theta(|V|d)$ term and lowers $\tau_{D}$. Furthermore, if this subset is not context-aware, the acceptance rate $\alpha$ may also drop, shrinking $\Omega(\alpha,\gamma)$ and hence offsetting the latency gains of vocabulary truncation.
4 DynaSpec - Context-aware Dynamic Sampling
In a nutshell, DynaSpec is a context-aware dynamic shortlisting method for speculative decoding with large-vocabulary LLMs, targeting the drafter’s output-head bottleneck that scales as $\mathcal{O}(|V|d)$. Instead of drafting over the full vocabulary (or a brittle, frequency-ranked static shortlist), DynaSpec uses a lightweight router/meta-classifier to select a small set of coarse token clusters for the current context, and the drafter evaluates logits only on the union of tokens in those clusters. Crucially, the target model still verifies over the full vocabulary, so decoding remains exact, while drafting becomes cheaper and better preserves acceptance on rare/domain tokens than fixed shortlists. Systems-wise, the router runs in parallel with draft encoding and DynaSpec uses a position-aware cluster budget (larger early, smaller later) with a fused indexed-head kernel to reduce overhead where it matters most.
Before presenting the training pipeline of DynaSpec, we setup the notation and outline (i) Eagle  (Li et al., 2024b, a) and (ii) FR-Spec Zhao et al. (2025), which are most relevant to our work.
4.1 Notation and preliminaries